OpenStudy (anonymous):

how do I find G(x)= 3x + IxI over x ?

OpenStudy (anonymous):

\[G(x)=\frac{x+|x|}{x}\] depends on whether x is positive or negative. if x is negative then it is 0

OpenStudy (anonymous):

if x is positive it is 2 and if x is zero it is undefined

OpenStudy (anonymous):

ok you had \[G(x)=\frac{3x+|x|}{x}\] make the necessary changes and you will have it

OpenStudy (anonymous):

how did u get two for the answer or is that a rule ?

OpenStudy (anonymous):

what is always true is that \[|x|= x \text{ if } x\geq 0\] and \[|x|=-x \text{ if } x < 0\] so you have to break it up into two cases i guess you could say it is a rule; it is the definition of absolute value

OpenStudy (anonymous):

easy with an example. suppose x was -2 then |x|=2 and you would have \[\frac{3\times -2+2}{-2}=\frac{-6-2}{-2}=\frac{-8}{-2}=2\] any other negative number will give you two as well. try it

OpenStudy (anonymous):

aah yes, can u explain why my book answer is \[(-\infty,0) \cup (0,\infty)\]